What is the probability that two female students will be selected at random, one after the other, to participate in a research project from a class of 7 male and 3 female students?

a. 0.0667
b. 0.3
c. 0.1
d. 0.2222
1
Expert's answer
2020-12-28T14:56:36-0500

The total number of people is 10.

The number of selecting two female students is a combination:

"C(3,2)=\frac{3!}{2!1!}=3"

The number of all possible outcomes selecting two people:

"C(10, 2)=\frac{10!}{2!8!}=45"

Hence the probability of selecting 2 female students is "\frac{3}{45}=0.0667"


Answer: a